Part A. Organize the data Drag the progrebots to the resectie targets View
Answer #1
analyze the problem initial condition final condition known predict
pressure 20 atm 3.0 atm decreases
volume 1.0 L increases

Now, pressure is inversely proportional to volume at constant temperature

so, P \propto\frac{1}{V}

or, P1V1 = P2V2

or, 20\times 1.0 = 3.0 \times V2

or, V2 = 6.67 L

hence final volume of helium = 6.67 L.